# Multiple Imputation for Nonresponse in Complex Surveys Using Design Weights and Auxiliary Margins

### v3

@@ 1 Introduction
In this short note, we consider settings where analysts seek to integrate auxiliary information on marginal distributions in multiple imputation for nonresponse (Rubin, 1987) in complex surveys. We build on the missing data with auxiliary margins, or MD-AM, framework introduced by (Akande et al., 2021) and extended to complex surveys by Akande and Reiter (2022), (Tang et al., 2024), and (Yang and Reiter, 2025). The latter three works impute missing values to ensure completed-data, survey-weighted inferences are plausible given the known margins. However, these works impose a simplifying condition on the unit nonrespondents, namely that the analyst does not use their design weights and instead replaces these weights with a convenient constant. This creates difficulties for design-based analyses, as the resulting weights for the full sample do not correspond to the design. Further, treating all unit nonrespondents exchangeably discards information that could inform imputations.
We propose MD-AM models that allow use of the design-based weights for all sampled units. The basic idea is as follows. We first impute all missing values due to item nonresponse using multiple imputation by chained equations (Van Buuren, 2018; Raghunathan et al., 2001). In each partially completed dataset, we then impute the unit nonrespondents’ values of the categorical variables that have auxiliary margins. To do so, we specify initial probabilities for imputations based on logistic regressions estimated using the completed data. We adjust these probabilities so that, in expectation, imputations result in plausible Horvitz and Thompson (1952) estimates in the completed data given the known margins, using the design weights for all sampled units. Finally, for the variables without auxiliary margins, we use the random hot deck imputation proposed in Yang and Reiter (2025).
The remainder of this article is organized as follows. Section 2 provides a brief review of the hybrid missing MD-AM model (Tang et al., 2024; Yang and Reiter, 2025), which we extend to incorporate design weights for unit nonrespondents. Section 3 introduces the methodology. Section 4 illustrates the methods using simulation studies. Code and additional simulation results are available at URL HERE.
@@ 2 Hybrid Missingness MD-AM Modeling
Let $U$ be the unit nonresponse indicator so that, for each unit $i\in\mathcal{S}$, $U_{i}=1$ if the unit provides no responses to any survey questions and $U_{i}=0$ otherwise. Let $\boldsymbol{R}=(R_{1},\dots,R_{p})$ be item nonresponse indicators corresponding to $(X_{1},\dots,X_{p})$ so that, for any unit $i$ with $U_{i}=0$ and any variable $X_{j}$, $R_{ij}=1$ if unit $i$ does not respond to question $X_{j}$ and $R_{ij}=0$ otherwise. When $U_{i}=1$, $(R_{i1},\dots,R_{ij})$ are undefined.
The survey organization has auxiliary information $\mathcal{A}$ on the marginal distributions of a subset of the $p$ survey variables. For $j=1,\dots,p$, let $\mathcal{A}_{j}$ comprise the known marginal in $\mathcal{P}$ for $X_{j}$, when it exists. We write $X_{j}\in\mathcal{A}$ whenever $X_{j}$ has a known margin in $\mathcal{A}$. For convenience, we presume each $\mathcal{A}_{j}$ comprises totals; incorporating percentages is a trivial modification. For any categorical $X_{j}$ taking on levels $c=\{1,\dots,m_{j}\}$ and $X_{j}\in\mathcal{A}$, let $T_{jc}=\sum_{i=1}^{N}I(x_{ij}=c)$ be the total number of units in $\mathcal{P}$ at level $c$. Here, $I(\cdot)=1$ when the condition inside the parenthesis is true and $I(\cdot)=0$ otherwise.
@@
$w_{i}=\left\{\begin{array}[]{lr}w_{i}^{d}&\text{if }U_{i}=0\\ \frac{N-\sum_{j\in\mathcal{S}}w_{j}^{d}}{\sum_{j\in\mathcal{S}}U_{j}}&\text{if}U_{i}=1.\end{array}\right.$
For any unit $i\in\mathcal{S}$ and for all $j$, let $x_{ij}^{\star}=x_{ij}$ when $(R_{ij}=0,U_{i}=0)$, and let $x_{ij}^{\star}$ be an imputed value when $R_{ij}=1$ or $U_{i}=1$. For any $X_{j}\in\mathcal{A}$, Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025) suppose the imputations to adhere to (2.1) – (2.3), with the additional constraint that
@@ 3.2 Variables with Margins for Unit Nonrespondents
Let $p_{i2c}=\Pr(X_{i2}=c|x_{i1}^{*}=d,w_{i})$ be the working probability of $X_{2}=c$ for unit $i$, given its imputed $x_{i1}^{*}=d$ and $w_{i}$. We determine these from a (multinomial) logistic regression of $X_{2}$ on $X_{1}$ and possibly some function of $W$, with coefficients estimated from the completed data for units with $U_{i}=0$. If we disregard $W$, $p_{i2c}=\Pr(X_{2}=c|X_{1}=d)$ for $c=1,\dots,m_{2}$ and $d=1,\dots,m_{1}$; that is, the conditional probability is the same for all units with a common value of $x_{i1}^{*}=d$.
@@
We impute each $x_{i2}^{*}$ given $x_{i1}^{*}$ for units with $U_{i}=1$ using a random draw from the adjusted probability mass function $\{\tilde{p}_{i2c}:c=1,\dots,m_{2}\}$, where $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}$ for $c=1,\dots,m_{2}-1$ and $\tilde{p}_{i2m_{2}}=1-\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}$. If $\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}>1$, we compute (3.3) for $c=1,\dots,m_{2}$ and set $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}/\sum_{c=1}^{m_{2}}f_{2c}{p}_{i2c}$. This process ensures the completed-data Horvitz and Thompson (1952) estimators approximately match the sampled $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}})$ in expectation, while also incorporating relationships between $X_{2}$ and $(X_{1},W)$ implied in the working probabilities. We independently repeat the estimation and imputation process in each completed dataset.
@@
When disregarding $W$ in the logistic regression for $X_{2}$ on $X_{1}$—so that within any completed dataset, $p_{i2c}=\Pr(X_{2}=c|x_{1}=d,U=1)$ for all unit nonrespondents with $x_{i1}^{*}=d$ in that dataset—we can find $\{\tilde{p}_{2c}\}$ by solving a system of equations implied by (2.2). Specifically, we encode the constant log-odds assumption as a set of $(m_{1}-1)(m_{2}-1)$ equations. For $c=2,\dots,m_{2}-1$ and $d=2,\dots,m_{1}-1$, we have
$\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=1)}{\Pr(X_{2}=1|X_{1}=d,U=1)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=1)}{\Pr(X_{2}=c|X_{1}=1,U=1)}\right)$
$=\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=0)}{\Pr(X_{2}=1|X_{1}=d,U=0)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=0)}{\Pr(X_{2}=c|X_{1}=1,U=0)}\right).$ (3.4)
We estimate the conditional probabilities for units with $U_{i}=0$ via survey-weighted ratio estimators using the completed data; for example,
@@
To illustrate, we describe the computations for (3.4) and (3.6) for binary $X_{2}$ and $X_{1}$. We use these in the simulations of Section 4. First, we compute $\Pr(X_{2}=1|X_{1}=1,U=0)$ and $\Pr(X_{2}=1|X_{1}=0,U=0)$ using (3.5). To ease notation, define
@@ 4 Simulation Studies
To demonstrate the methodology, we mimic the simulation study in Yang and Reiter (2025). They found that, by leveraging auxiliary margins, multiple imputation via the hybrid missingness MD-AM model can offer more accurate inferences than multiple imputation that does not utilize $\mathcal{A}$. However, they presume design weights are unused or unavailable for unit nonrespondents, whereas we use them per Section 3.
We generate $\mathcal{P}$ comprising $N=3,373,378$ units. This is the number of individuals in the 2022 American Community Survey data file used by Yang and Reiter (2025). We make one of the variables in this file a size variable $Z=(z_{1},\dots,z_{N})$; this is observed for all $N$ individuals. To generate a $\mathcal{S}$, we sample each record from $\mathcal{P}$ independently with $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling design results in approximately $n=6000$ units in any $\mathcal{S}$. Let $W=10Z$ be the survey weights. For each unit $i=1,\dots,N$, we generate four binary variables $(x_{i1},x_{i2},x_{i3},x_{i4})$ and two continuous variables $(x_{i5},x_{i6})$. The data generation model is displayed in (4.1)–(4.5).
@@
We independently sample 500 datasets. For each sample $\mathcal{S}$, we generate the unit nonrespondents by sampling Bernoulli random draws of $U_{i}$, where $i\in\mathcal{S}$, from the distribution of $U|X_{1},X_{2}$ implied by (4.1)–(4.3). We make $\boldsymbol{x}_{i}$ completely missing for any unit in $\mathcal{S}$ where the drawn $U_{i}=1$. For item nonresponse, we use the ICIN mechanisms defined in (4.6) for $X_{j}$ where $j\in\{2,3,4,6\}$, and we let $X_{1}$ and $X_{5}$ be fully observed.
$R_{j}|X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},U=0\sim\text{Bernoulli}(\pi_{R_{j}}),\,\,\,\text{logit}(\pi_{R_{j}})=\phi_{j0}+\sum_{k\neq j}\phi_{jk}X_{k}.$ (4.6)
@@
We blank every $x_{ij}$ for which the sampled $R_{ij}=1$. We set the parameters in (4.6) so that approximately 25% of each eligible survey variable is missing for units with $U_{i}=0$. Specific parameter values are available in Appendix A.
@@
We let $\mathcal{A}$ include $T_{1}=\sum_{i=1}^{N}x_{i1}$ and $T_{2}=\sum_{i=1}^{N}x_{i2}$, which we use for multiple imputation via the MD-AM models of Section 3. For Section 3.1, we generate $L=10$ completed datasets for item nonresponse using a bespoke implementation of the “mice” package in R. For Section 3.2, we consider two sets of working probabilities for $X_{1}$. The first allows $p_{i1c}$ to vary across units: each $p_{i1c}$ equals the unit’s predicted probability from the logistic regression of $X_{1}$ on $W$ estimated using the units with $U_{i}=0$. The second fixes $p_{i1c}=p_{1c}$ across units and uses (3.2). For both we use (3.9) and (3.10) to impute $X_{2}$. For the variances $V_{1}$ and $V_{2}$ used to sample $\hat{T}_{1}\sim N(T_{1},V_{1})$ and $\hat{T}_{2}\sim N(T_{2},V_{2})$, we use “mice” to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the unbiased variance estimators for $\hat{T}_{1}$ and $\hat{T}_{2}$ under Poisson sampling. For Section 3.3, we create donor pools for the hot deck by matching on $(X_{1},X_{2})$.
@@
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the estimands in Yang and Reiter (2025), including marginal totals for $(X_{1},\dots,X_{6})$ and several probabilities involving multiple categorical variables. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators for a Poisson sampling design based on the design weights for all sampled records.
@@
Figure XX summarizes the results for $\theta_{1}=-2$; results for $\theta_{1}=-0.5$ are available at URL HERE. Both MD-AM models produce accurate point estimates, mitigating the substantial bias from the nonignorable nonresponse. The multiple imputation confidence intervals have coverage rates exceeding 95% due to over-estimation of the variances. As discussed by Yang and Reiter (2025), this is because the combining rules of Rubin (1987) do not account for using known totals in imputation. Overall, the simulation results suggest that the MD-AM models from Section 3 offer principled ways to incorporate known marginal totals in multiple imputation when using the design weights for all sampled units.
Appendix A Simulation Study Parameters

### v4

@@ 1 Introduction
In this short note, we consider settings where analysts seek to integrate auxiliary information on marginal distributions in multiple imputation for nonresponse (Rubin, 1987) in complex surveys. We build on the missing data with auxiliary margins, or MD-AM, framework introduced by Akande et al. (2021) and extended to complex surveys by Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025). The latter three works impute missing values to ensure completed-data, survey-weighted inferences are plausible given the known margins. However, these works impose a simplifying condition on the unit nonrespondents, namely that the analyst does not use their design weights and instead replaces these weights with a convenient constant. This creates difficulties for design-based analyses, as the resulting weights for the full sample do not correspond to the design.
We propose MD-AM models that allow use of the design-based weights for all sampled units. The basic idea is as follows. We first impute all missing values due to item nonresponse using multiple imputation by chained equations (Van Buuren, 2018; Raghunathan et al., 2001). In each partially completed dataset, we then impute the unit nonrespondents’ values of the categorical variables that have auxiliary margins. To do so, we specify probabilities for imputations so that, in expectation, imputations result in plausible Horvitz and Thompson (1952) estimates in the completed data given the known margins, using the design weights for all sampled units. For the variables without auxiliary margins, we use the random hot deck imputation proposed in Yang and Reiter (2025).
The remainder of this note is organized as follows. Section 2 provides a brief review of the hybrid missing MD-AM model (Tang et al., 2024; Yang and Reiter, 2025), which we extend to incorporate design weights for unit nonrespondents. Section 3 introduces the methodology. Section 4 illustrates the methods using simulation studies. Codes are available at https://github.com/kevinxu47/MDAM.
@@ 2 Hybrid Missingness MD-AM Modeling
Let $U$ be the unit nonresponse indicator so that, for each unit $i\in\mathcal{S}$, $U_{i}=1$ if the unit provides no responses to any survey questions and $U_{i}=0$ otherwise. Let $\boldsymbol{R}=(R_{1},\dots,R_{p})$ be item nonresponse indicators corresponding to $\boldsymbol{X}=(X_{1},\dots,X_{p})$ so that, for any unit $i$ with $U_{i}=0$ and any variable $X_{j}$, $R_{ij}=1$ if unit $i$ does not respond to question $X_{j}$ and $R_{ij}=0$ otherwise. When $U_{i}=1$, $(R_{i1},\dots,R_{ij})$ are undefined.
The survey organization has auxiliary information $\mathcal{A}$ on the marginal distributions of a subset of the $p$ survey variables. For $j=1,\dots,p$, let $\mathcal{A}_{j}$ comprise the known marginal for $X_{j}$ when it exists. We write $X_{j}\in\mathcal{A}$ whenever $X_{j}$ has a known margin in $\mathcal{A}$. For convenience, we presume each $\mathcal{A}_{j}$ comprises population totals; incorporating percentages is a simple modification. For any categorical $X_{j}$ taking on levels $c=\{1,\dots,m_{j}\}$ and $X_{j}\in\mathcal{A}$, let $T_{jc}=\sum_{i=1}^{N}I(x_{ij}=c)$ be the total number of units in $\mathcal{P}$ at level $c$. Here, $I(\cdot)=1$ when the condition inside the parenthesis is true and $I(\cdot)=0$ otherwise.
@@
$w_{i}=\left\{\begin{array}[]{lr}w_{i}^{d}&\text{if }U_{i}=0\\ \frac{N-\sum_{k\in\mathcal{S}}w_{k}^{d}}{\sum_{k\in\mathcal{S}}U_{k}}&\text{if}U_{i}=1.\end{array}\right.$ (2.4)
For any unit $i\in\mathcal{S}$ and $j=1,\dots,p$, let $x_{ij}^{\star}=x_{ij}$ when $(R_{ij}=0,U_{i}=0)$, and let $x_{ij}^{\star}$ be an imputed value when $R_{ij}=1$ or $U_{i}=1$. For any $X_{j}\in\mathcal{A}$, Akande and Reiter (2022), Tang et al. (2024), and Yang and Reiter (2025) suppose the imputations to adhere to (2.1) – (2.3), with the additional constraint that
@@ 3.2 Variables with Margins for Unit Nonrespondents
Let $p_{i2c}=\Pr(X_{i2}=c|x_{i1}^{*}=d,w_{i})$ be the working probability that $X_{2}=c$ for unit $i$, given its imputed $x_{i1}^{*}=d$ and $w_{i}$. We determine these from a (multinomial) logistic regression of $X_{2}$ on $X_{1}$ and possibly some function of $W$, with coefficients estimated from the completed data for units with $U_{i}=0$. If we disregard $W$, $p_{i2c}=\Pr(X_{2}=c|X_{1}=d)$ for $c=1,\dots,m_{2}$ and $d=1,\dots,m_{1}$; that is, the conditional probability is the same for all units with a common value of $x_{i1}^{*}=d$.
@@
We impute each $x_{i2}^{*}$ given $x_{i1}^{*}$ for units with $U_{i}=1$ using a random draw from the adjusted probability mass function $\{\tilde{p}_{i2c}:c=1,\dots,m_{2}\}$, where $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}$ for $c=1,\dots,m_{2}-1$ and $\tilde{p}_{i2m_{2}}=1-\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}$. If $\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}>1$, we compute (3.3) for $c=1,\dots,m_{2}$ and set $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}/\sum_{c=1}^{m_{2}}f_{2c}{p}_{i2c}$. This process ensures the completed-data Horvitz and Thompson (1952) estimates approximately match the sampled $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}})$ in expectation, while also incorporating relationships between $X_{2}$ and $(X_{1},W)$ implied in the working probabilities. We independently repeat the estimation and imputation process in each completed dataset.
@@
When disregarding $W$ in the logistic regression for $X_{2}$ on $X_{1}$—so that within any completed dataset, $p_{i2c}=\Pr(X_{2}=c|x_{1}=d,U=1)$ for all unit nonrespondents with $x_{i1}^{*}=d$ in that dataset—we can find $\{\tilde{p}_{2c}\}$ by solving a system of equations implied by (2.2). Specifically, we encode the constant log-odds assumption as a set of $(m_{1}-1)(m_{2}-1)$ equations. For $c=2,\dots,m_{2}$ and $d=2,\dots,m_{1}$, we have
$\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=1)}{\Pr(X_{2}=1|X_{1}=d,U=1)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=1,U=1)}{\Pr(X_{2}=1|X_{1}=1,U=1)}\right)$
$=\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=0)}{\Pr(X_{2}=1|X_{1}=d,U=0)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=1,U=0)}{\Pr(X_{2}=1|X_{1}=1,U=0)}\right).$ (3.4)
We can estimate the conditional probabilities for units with $U_{i}=0$ via survey-weighted ratio estimators using the completed data; for example,
@@
To illustrate, we present the computations for (3.4) and (3.6) for binary $X_{2}$ and $X_{1}$. We use these in the simulations of Section 4. First, we compute $\Pr(X_{2}=1|X_{1}=1,U=0)$ and $\Pr(X_{2}=1|X_{1}=0,U=0)$ using (3.5). To ease notation, define
@@ 4 Simulation Studies
To demonstrate the methodology, we mimic the simulation study in Yang and Reiter (2025). They found that, by leveraging auxiliary margins, multiple imputation via the hybrid missingness MD-AM model can offer more accurate inferences than multiple imputation that does not utilize $\mathcal{A}$. However, they use (2.4) in lieu of the design weights for unit nonrespondents, whereas we use them per Section 3.
We generate $\mathcal{P}$ comprising $N=3,373,378$ units. This is the number of individuals in the 2022 American Community Survey data file used by Yang and Reiter (2025). We make one of the variables in this file a size variable $Z=(z_{1},\dots,z_{N})$; this is observed for all $N$ individuals. To generate a $\mathcal{S}$, we sample records from $\mathcal{P}$ independently with $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling design results in approximately $n=6000$ units in any $\mathcal{S}$. Let $W=10Z$ be the survey weights.
For each unit $i=1,\dots,N$, we generate four binary variables $(x_{i1},x_{i2},x_{i3},x_{i4})$ and two continuous variables $(x_{i5},x_{i6})$. The data generation model is displayed in (4.1)–(4.5).
@@
We independently sample 500 datasets. For each sample $\mathcal{S}$, we generate the unit nonrespondents by sampling Bernoulli random draws of $U_{i}$, where $i\in\mathcal{S}$, from the distribution of $U|X_{1},X_{2}$ implied by (4.1)–(4.3). We make $\boldsymbol{x}_{i}$ completely missing for any unit in $\mathcal{S}$ where the drawn $U_{i}=1$. For item nonresponse, we use the ICIN mechanisms
$R_{j}|X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},U=0\sim\text{Bernoulli}(\pi_{R_{j}}),\,\,\,\text{logit}(\pi_{R_{j}})=\phi_{j0}+\sum_{k\neq j}\phi_{jk}X_{k}$ (4.6)
@@
for $X_{j}$ where $j\in\{2,3,4,6\}$, and we let $X_{1}$ and $X_{5}$ be fully observed. We blank every $x_{ij}$ for which the sampled $R_{ij}=1$. We set parameters in (4.6) so that approximately 25% of each eligible survey variable is missing for units with $U_{i}=0$. Specific parameter values are available in Appendix A.
@@
We let $\mathcal{A}$ include $T_{1}=\sum_{i=1}^{N}x_{i1}$ and $T_{2}=\sum_{i=1}^{N}x_{i2}$, which we use for multiple imputation via the MD-AM models of Section 3. Using Section 3.1, we generate $L=10$ completed datasets for item nonresponse using a bespoke implementation of the “mice” package in R. Using Section 3.2, we consider two sets of working probabilities for $X_{1}$. The first allows $p_{i1c}$ to vary across units: each $p_{i1c}$ equals the unit’s predicted probability from the logistic regression of $X_{1}$ on $W$ estimated using the units with $U_{i}=0$. The second fixes $p_{i1c}=p_{1c}$ across units and uses (3.2). For both we use (3.9) and (3.10) to impute $X_{2}$. For the variances $V_{1}$ and $V_{2}$ used to sample $\hat{T}_{1}\sim N(T_{1},V_{1})$ and $\hat{T}_{2}\sim N(T_{2},V_{2})$, we use “mice” to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the unbiased variance estimators for $\hat{T}_{1}$ and $\hat{T}_{2}$ under Poisson sampling. Using Section 3.3, we create donor pools for the hot deck by matching on $(X_{1},X_{2})$.
@@
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the estimands in Yang and Reiter (2025), including marginal totals for $(X_{1},\dots,X_{6})$ and several probabilities involving multiple categorical variables. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators for a Poisson sampling design based on the design weights for all sampled records. For comparisons, we also use the imputation strategy in Yang and Reiter (2025) that does not utilize $\mathcal{A}$. It uses a bespoke implementation of “mice” to impute missing items and randomly samples whole records with replacement as imputations for unit nonresponse. We refer to this method as IH.
@@
Figure 4.1 summarizes the relative root mean squared errors (RMSEs) of the point estimates and empirical coverage rates of 95% multiple imputation confidence intervals for the three methods. Results in tabular form are at https://github.com/kevinxu47/MDAM. Both MD-AM models produce accurate point estimates, mitigating the bias from the nonignorable nonresponse evident in IH. The coverage rates for both MD-AM models tend to exceed 95% due to over-estimation of variances. As discussed by Yang and Reiter (2025), this is because the combining rules of Rubin (1987) do not account for using known totals in imputation. Overall, the results suggest that the MD-AM models enable analysts to incorporate known marginal totals in multiple imputation when using all sampled units’ design weights.
Figure 4.1: Simulated relative RMSEs and coverage rates for population totals, two-way conditional probabilities, and joint probabilities for MD-AM and IH models. Left panel compares IH to the MD-AM model with fixed working probabilities. Right panel displays the two MD-AM models. Top row uses $\theta_{1}=-2$. Bottom row uses $\theta_{1}=-0.5.$
Appendix A Simulation Parameters